Can someone help me solve this?
Type of Problem: Applications of Linear Systems
Solve the following problem.
The Nile River is 394 kilometers longer than the Amazon River. The two rivers have a total length of 12,986 kilometers. How long is each river?

Answer by Cromlix(4381)

Amazon River = x
Nile River = x + 394km
x + (x + 394) = 12,986km
x + x + 394 = 12,986km
2x = 12,986 - 394
2x = 12592km
x = 6296km
Amazon River = 6296km
Nile River = 6690km.
Hope this helps.
:-)